OpenStudy (anonymous):

The sum of the measures of a complement and a supplement of an angle is 200. What is the measure of the angle? A. 70 B. 55 C. 20 D. 35

OpenStudy (radar):

Lets call this angle x. Then 90-x is the compliment. Ergo, 180-x is the supplement. The problem states (90-x) + ((180-x)=200. Next step is to solve that equation. 90 -x + 180 - x=200 -2x + 270 = 200 -2x = -70 x = 35 The supplement of 35 is 145 The compliment of 35 is 55 and sure enought 145 + 55 = 200 answer thus verified.

OpenStudy (radar):

This problem is testing if you know the definitions of supplementary and complementary angles and if you can do a little algebra.